Solve the equation: -4x - 7y -3z = 14

Answer should be in this format:

5484844
Answer #1

Given - An-74-32= 14 -> - 4x = 14 + 7y +32 Let y=8 and a=t 3 un = 78+ 3t +14. *=ms +2+ +-14 y= s tot to. y = ott to 1-3/4 *B-